what is the quotient of 1 2/5 and 4 5/6
Answer by juaneleanor(3)

Convert mixed fractions to improper fractions

1 2/5=7/5

4 5/6=29/6

Now divide the 2 improper fractions

7/5÷29/6

Since its division, you must get the reciprocal of the divisor, and multiply the 2 fractions

7/5×6/29

Therefore the answer is…

42/145


Eleanor